# Dynamical disequilibrium in dwarf galaxies: rethinking gas dynamics, rotation curves, and dark matter inference

### v1

@@ 2.3 Dynamics of gas discs: fluid models & circular velocity reconstruction
Further simplifications arise by assuming an idealized, axisymmetric disc with negligible radial and vertical motions, and where the pressure gradient is either small or corrected separately (e.g. via ‘asymmetric drift corrections’; see Valenzuela2007, appendix A, for a helpful discussion of this technique and related terminology). Under these conditions, the radial convective term reduces to the centrifugal acceleration, yielding the familiar relation:
@@ 2.4 SPH-based continuous field sampling
This continuous-field reconstruction extends particle-based analyses into smooth, spatially resolved descriptions of the gas dynamics. Technically, it corresponds to a Shepard (Shepard1968) SPH–gather scheme employing the same kernel definitions and neighbour criteria as the simulation itself, thereby ensuring consistency with its hydrodynamical framework. An important advantage of this scheme is that each sampling point attains an effective spatial resolution determined self-consistently by the local fluid density. This guarantees that reconstructed quantities vary smoothly across the disc, mitigating the shot noise typical of scatter-type interpolations, where individual particle contributions dominate the local estimate. The Shepard normalization further enforces that the effective kernel sum satisfies $\sum_{i}\tilde{W}=1$, eliminating one of the main sources of bias in standard SPH interpolation – deviations from unity kernel normalization – which can otherwise lead to systematic under- or overestimation of field amplitudes in sparsely sampled regions.
A caveat of standard Shepard formalisms is that they are not strictly conservative in mass or momentum (Reinhardt2019). However, since the sampling points are not interpreted as finite-volume fluid cells, but as infinitesimal tracers defining a continuous warped midplane, the lack of strict conservation poses no conceptual inconsistency for the subsequent equilibrium analysis. Our reconstruction does not attempt to reproduce the exact numerical forces used during the simulation’s time integration, but rather to recover the instantaneous, continuous field structure implied by the discrete SPH representation.
When evaluating the hydrodynamical Equations (1)–(2), we adopt several deliberate departures from the native SPH formulation used during the simulation. In particular, we exclude (i) artificial viscosity terms, which are numerical constructs introduced for entropy dissipation and shock capturing (Borrow2022), and (ii) the contribution from the large-scale gravitational field of the full simulation box. Our focus is instead on the physical force balance of the gas under the assumption of an inviscid, collisional fluid.
@@ 3.1 Local performance of fluid models
First, all reconstructions exhibit significant azimuthal variability at fixed radius, indicating strong model-dependent sensitivity to local dynamics and asymmetries. Such non‑axisymmetric structures are well documented in empirical gas discs’ hydrodynamical fields (e.g., harmonic decomposition of moment maps revealing spiral patterns and/or n-symmetric velocity fields; see Schoenmakers1997; Trachternach2008; Sellwood2010). However, the behaviours observed here underscore a deeper implication: azimuthal averaging, which is commonly used in rotation curve analyses (both observational and theoretical), can obscure critical localised discrepancies. These deviations are physical signatures of disequilibrium physics, driven by ongoing astrophysical processes (as we will discuss in Sec. 4).
Second, and perhaps more strikingly, the ‘Equilibrium model’ (dark green/dotted) – despite being the most comprehensive physical model – is frequently outperformed by the simplistic $v_{\phi}=v_{c}$ assumption (orange/dashed); see, for example, the increased absolute deviation of the equilibrium reconstruction from the true circular velocity in the $\phi^{\prime}=180^{\circ}$ slice. The simpler ‘Pressure support model’ (brown/dash-dotted), which attempts a correction to $v_{\phi}$ under the assumption that hydrostatic pressure gradients are the dominant non-equilibrium effect, fares similarly – in this galaxy this assumption is justified across most of the disc, although this does not hold for all galaxies. This counterintuitive outcome echoes the findings of Sands2024 who studied disequilibrium in gas discs from the FIRE-3 simulation suite (Hopkins2023). They argue that the poor performance of the equilibrium model can be traced to the unbalanced nature of Eulerian terms. Specifically, radial components of the convective derivative, $(\boldsymbol{v}\cdot\nabla)\boldsymbol{v}$, often arise from the same non-equilibrium processes that generate time variations in the velocity field, $\partial_{t}\boldsymbol{v}$. This coupling undermines the validity of equilibrium reconstructions. That we find similar trends in COLIBRE reinforces a key point: adding correction terms for non‑circular motions or pressure support does not guarantee more accurate mass inferences. The limiting factor is whether the gas is actually in hydrodynamical equilibrium.
@@ 3.2 Population-wide radial trends in disequilibrium physics
Figure 2: Radial distributions of the $\varepsilon(r,\phi)$ and $\delta(r,\phi)$ parameters (Equation (4)) weighted by gas mass across the full galaxy sample, respectively $\mathrm{W}(\varepsilon(r,\phi),r)$ and $\mathrm{W}(\delta(r,\phi),r)$. Each panel displays both the joint and marginal distributions of the respective parameter. Left panel: Stacked 2D histogram of $\varepsilon$ vs. galactocentric radius, $r$, where the colour scale indicates the joint weighted density, $\mathrm{W}(\varepsilon,r)/\Sigma\mathrm{W}$, summed across all galaxies. The vertical subpanel to the right shows the normalised marginal (1D) distribution, $\mathrm{W}(\varepsilon)/\Sigma\mathrm{W}$. Right panel: Same as left panel, but for the $\delta$ parameter. Cyan dash-dotted lines in the main panel and black dash-dotted lines in the side panel denote the parameter range associated with $\leq 10$ per cent error in circular velocity inference based on the corresponding dynamical model.
@@
Fig. 2 shows the radial distribution of both diagnostic parameters across the galaxy sample. The left panel corresponds to the mass-weighted density of $\varepsilon(r,\phi)$ – i.e. $\mathrm{W}(\varepsilon,r)/\Sigma\mathrm{W}$ – while the right panel shows the same for $\delta(r,\phi)$. In each case, the main subpanel presents a stacked 2D histogram of parameter values as a function of galactocentric radius, with colour indicating the joint distribution normalised by the total counts. Each galaxy here contributes 180 evenly-spaced azimuthal measurements per radial bin up to the radius marking the extent of its gaseous disc – $R_{\rm ext,H\textsc{i}}$, as defined in Sec. 2.2, enabling robust population-level statistics. The vertical subpanels display the normalised 1D distribution of the corresponding parameter. Cyan (main panels) and black (side panels) dash-dotted lines indicate the range of parameter values associated with $\leq 10$ per cent error in circular velocity inference under the relevant fluid model (i.e. $|\varepsilon|\lesssim 0.2$, $|\delta|\leq 0.1$). Several results emerge.
@@
As evident in the main panels, the vast majority of the samples contributing to these broad distributions originate from the inner disc regions ($r\lesssim 5$ kpc). This finding runs counter to expectations based on dynamical timescales, which predict more rapid relaxation in the central regions; instead, we find clear evidence for persistent perturbations. This is particularly relevant for core–cusp studies, since this radial range provides the strongest leverage for constraining dark matter density profiles. This is not the only hint that the centres of dwarf galaxies are continuously perturbed out of equilibrium. In fact, a recent investigation of radial trends in the radial acceleration scaling relation (RAR; McGaugh2016) using the SPARC sample (Lelli2016) shows that the inner regions – defined as $r<2R_{d}$, where $R_{d}$ is the stellar disc scale radius – are much less tight (in terms of intrinsic scatter) than the outer disc (see Fig. S1 in Ren2019). This trend can be interpreted as empirical evidence that inner regions of galaxies are indeed the most dynamically disturbed.
@@
Moreover, the $\delta$ parameter reveals a population-wide tendency to underestimate the true circular velocity in the inner disc ($\delta<0$), implying a systematic bias toward inferring smaller inner velocities, hence ‘cored’ profiles. However, the distribution is broad, with a minority of regions exhibiting the opposite trend ($\delta>0$). This bidirectional scatter suggests that some of the observed diversity in dwarf galaxy rotation curves could arise from disequilibrium physics rather than intrinsic variations in the underlying dark matter distributions. In other words, the rotational structure in the inner regions of gas discs is a biased and high-variance tracer of the underlying gravitational potential even before projection effects, and the limited ability of kinematic models to accurately recover azimuthal speeds, come into play.
@@ 4.1.1 Stellar feedback
Stellar feedback is a well-motivated source of ISM disequilibrium, particularly in shallow gravitational potentials where momentum and energy injection from massive stars can rival the disc binding energy (e.g. Dekel1986; Stinson2006; Hopkins2012; KimOstriker2017). Energy deposition from SNe, stellar winds, and radiation pressure drives expanding shells, turbulence, and large-scale pressure gradients, each of which can displace gas from local force balance for timescales exceeding the local dynamical time (Martizzi2015; Orr2020). Observations support this picture: nearby dwarfs often exhibit asymmetric Hi profiles (e.g. Walter2008; Stilp2013) and vertically extended warm gas layers (Boomsma2008; Bacchini2020), with significant dispersive motions consistent with feedback-driven disturbances. A striking observational example of extreme disruption is the supergiant Hi shell, or ‘superbubble’, in IC2574 (Brinks2003), which also coincides with one of the most pronounced apparent dark matter cores known in a dwarf galaxy.
In our analysis, feedback-affected gas is defined as material lying within a projected radius of $1~\mathrm{kpc}$ in the warped disc midplane around either young stars (ages $\tau_{\star}\lesssim 100~\mathrm{Myr}$) or gas particles that have recently received kinetic injection from CCSNe ($t_{\rm lb}\leq 50~\mathrm{Myr}$). Particle positions are first assigned to their nearest point on the deprojected disc surface, and the selection is then made using a two-dimensional circular mask of radius $1~\mathrm{kpc}$ (see Sec. 2.5 for details). This choice reflects both the geometry of the discs and the physical dispersion of feedback-driven material: although CCSNe impart an isotropic target kick of $v_{\rm kick}=50~\mathrm{km~s^{-1}}$ in the rest frame of the star particle (Chaikin2023; Schaye2025), the residual velocity of affected gas in the rotating disc frame is typically smaller ($\sim 20~\mathrm{km~s^{-1}}$), implying characteristic displacements of $\sim 1~\mathrm{kpc}$ over $\sim 50~\mathrm{Myr}$ (ignoring other forces). Furthermore, Hi-dominated dwarf galaxy discs have expected vertical scale heights of $\sim 1~\mathrm{kpc}$ (Banerjee2011), making this spatial cut a physically motivated and conservative choice for identifying feedback-influenced regions in a population-wide analysis.
@@
Statistically (right panel), the stellar feedback-affected gas exhibits a strongly skewed $\varepsilon$ distribution with a pronounced tail towards large negative values, indicating a time-dependent, Eulerian, radially inward acceleration that offsets the excess outward force imparted by feedback. Physical intuition for this behaviour can be developed by first considering the rest frame of the young star launching the feedback. In this frame, a supernova injects an approximately isotropic impulse, producing a transient increase in the local radial velocity that decays as the blast wave evolves (Sedov1946; Taylor1950; VonNeumann1964); i.e. $\partial_{t}v_{r}^{\prime}<0$, where $v_{r}^{\prime}$ is the radial component measured away from the star. When transformed to the galactocentric frame, one might expect the inward- and outward-directed components of this symmetric impulse to contribute equally to $\partial_{t}v_{r}$ at a fixed location in the disc. However, the strong radial and vertical density gradients in galactic gas discs make it easier for feedback to displace gas outward or out of the plane of the disc than inward. As a result, the Eulerian $\partial_{t}v_{r}$ in the galactocentric frame is systematically biased negative when averaged across galaxies and mass-weighted, despite the underlying isotropy of the impulse.
@@ 4.1.2 Overdense and self-gravitating gas clumps
Local gravitational collapse within discs provides a natural route to hydrodynamical disequilibrium. Dense, self-gravitating clumps – often arising from gravitational instability in the cold ISM (e.g. Noguchi1999; Dekel2009; Genzel2011; Goldbaum2015) – develop internal dynamics that strongly deviate from the background shear flow.
@@
The middle row of Fig. 3 illustrates their spatial and dynamical imprint. In the left panel, the clump tracer field delineates spatially localised knots and filaments, frequently embedded within spiral arms or ring-like overdensities. The corresponding $\varepsilon$ map (middle) reveals that these same sites are often among the most extreme disequilibrium regions, with $\varepsilon$ commonly positive and exceeding unity. The right panel quantifies this: the clump-tagged distribution is markedly broader than the global disc-wide one, with a mild shift of the peak toward $\varepsilon\gtrsim 0$ and extended wings on both sides, reflecting a mix of accelerated collapse ($\varepsilon<0$) and migration ($\varepsilon>0$). This contrasts with the distribution characteristic of feedback zones, typically biased more strongly on the negative side (Sec. 4.1.1). In idealised scenarios, a clump drives a locally symmetric, converging flow in its own rest frame, such that $\partial_{t}v_{r}^{\prime}<0$, where $v_{r}^{\prime}$ is the radial component measured away from the clump (e.g., Penston1969). In the Eulerian galactocentric frame, because the ambient gas density decreases with radius, flux from the denser inward side exceed those from the outward side, biasing $\partial_{t}v_{r}$ towards the positive end. This naturally produces the mild positive shift in the clump-selected distribution. At the same time, many dense clumps either form within, or evolve into, regions where young stars subsequently inject feedback; sampling these mixed environments reintroduces the strongly negative tail characteristic of feedback-driven gas. The joint effect is a broad, wing-dominated distribution with a slight offset toward $\varepsilon>0$, as seen in Fig. 3.
While some clumps form or reside within feedback-active environments, a substantial fraction lie in quiescent or outer disc regions lacking obvious stellar-driven perturbations. This is evident in Fig. 3: because the same representative galaxy appears in the first two rows, the stellar-feedback (top-row) and clump-selected (middle-row) contours can be directly compared. Many overdense regions fall well outside the feedback-affected zones, yet still exhibit extreme $|\varepsilon|$, demonstrating that clump-driven dynamics generate strong disequilibria independently of stellar feedback. Observational analogues – such as the massive clumps in $z\sim 2$ discs (e.g. Elmegreen2007; Genzel2011), or in the nearby WLM dwarf galaxy (Albers2019; Mondal2021) – likewise show that such structures can persist without contemporaneous feedback. We therefore treat self-gravitating clumps as a separate, dynamically distinct driver of local disequilibrium.

### v2

@@ 2.3 Dynamics of gas discs: fluid models & circular velocity reconstruction
Further simplifications arise by assuming an idealized, axisymmetric disc with negligible radial and vertical motions, and where the pressure gradient is either small or corrected separately (e.g. via ‘asymmetric drift corrections’; see Valenzuela et al. 2007, appendix A, for a helpful discussion of this technique and related terminology). Under these conditions, the radial convective term reduces to the centrifugal acceleration, yielding the familiar relation:
@@ 2.4 SPH-based continuous field sampling
This continuous-field reconstruction extends particle-based analyses into smooth, spatially resolved descriptions of the gas dynamics. Technically, it corresponds to a Shepard (Shepard, 1968) SPH–gather scheme employing the same kernel definitions and neighbour criteria as the simulation itself, thereby ensuring consistency with its hydrodynamical framework. An important advantage of this scheme is that each sampling point attains an effective spatial resolution determined self-consistently by the local fluid density. This guarantees that reconstructed quantities vary smoothly across the disc, mitigating the shot noise typical of scatter-type interpolations, where individual particle contributions dominate the local estimate. The Shepard normalization further enforces that the effective kernel sum satisfies $\sum_{i}\tilde{W}=1$, eliminating one of the main sources of bias in standard SPH interpolation – deviations from unity kernel normalization – which can otherwise lead to systematic under- or overestimation of field amplitudes in sparsely sampled regions.
A caveat of standard Shepard formalisms is that they are not strictly conservative in mass or momentum (Reinhardt et al., 2019). However, since the sampling points are not interpreted as finite-volume fluid cells, but as infinitesimal tracers defining a continuous warped midplane, the lack of strict conservation poses no conceptual inconsistency for the subsequent equilibrium analysis. Our reconstruction does not attempt to reproduce the exact numerical forces used during the simulation’s time integration, but rather to recover the instantaneous, continuous field structure implied by the discrete SPH representation.
When evaluating the hydrodynamical Equations (1)–(2), we adopt several deliberate departures from the native SPH formulation used during the simulation. In particular, we exclude (i) artificial viscosity terms, which are numerical constructs introduced for entropy dissipation and shock capturing (Borrow et al., 2022), and (ii) the contribution from the large-scale gravitational field of the full simulation box. Our focus is instead on the physical force balance of the gas under the assumption of an inviscid, collisional fluid.
@@ 3.1 Local performance of fluid models
First, all reconstructions exhibit significant azimuthal variability at fixed radius, indicating strong model-dependent sensitivity to local dynamics and asymmetries. Such non‑axisymmetric structures are well documented in empirical gas discs’ hydrodynamical fields (e.g., harmonic decomposition of moment maps revealing spiral patterns and/or n-symmetric velocity fields; see Schoenmakers et al. 1997; Trachternach et al. 2008; Sellwood and Sánchez 2010). However, the behaviours observed here underscore a deeper implication: azimuthal averaging, which is commonly used in rotation curve analyses (both observational and theoretical), can obscure critical localised discrepancies. These deviations are physical signatures of disequilibrium physics, driven by ongoing astrophysical processes (as we will discuss in Sec. 4).
Second, and perhaps more strikingly, the ‘Equilibrium model’ (dark green/dotted) – despite being the most comprehensive physical model – is frequently outperformed by the simplistic $v_{\phi}=v_{c}$ assumption (orange/dashed); see, for example, the increased absolute deviation of the equilibrium reconstruction from the true circular velocity in the $\phi^{\prime}=180^{\circ}$ slice. The simpler ‘Pressure support model’ (brown/dash-dotted), which attempts a correction to $v_{\phi}$ under the assumption that hydrostatic pressure gradients are the dominant non-equilibrium effect, fares similarly – in this galaxy this assumption is justified across most of the disc, although this does not hold for all galaxies. This counterintuitive outcome echoes the findings of Sands et al. (2024) who studied disequilibrium in gas discs from the FIRE-3 simulation suite (Hopkins et al., 2023). They argue that the poor performance of the equilibrium model can be traced to the unbalanced nature of Eulerian terms. Specifically, radial components of the convective derivative, $(\boldsymbol{v}\cdot\nabla)\boldsymbol{v}$, often arise from the same non-equilibrium processes that generate time variations in the velocity field, $\partial_{t}\boldsymbol{v}$. This coupling undermines the validity of equilibrium reconstructions. That we find similar trends in COLIBRE reinforces a key point: adding correction terms for non‑circular motions or pressure support does not guarantee more accurate mass inferences. The limiting factor is whether the gas is actually in hydrodynamical equilibrium.
@@ 3.2 Population-wide radial trends in disequilibrium physics
Figure 2: Radial distributions of the $\varepsilon(r,\phi)$ and $\delta(r,\phi)$ parameters (Equation (4)) weighted by gas mass across the full galaxy sample, respectively $\mathrm{W}(\varepsilon(r,\phi),r)$ and $\mathrm{W}(\delta(r,\phi),r)$. Each panel displays both the joint and marginal distributions of the respective parameter. Left panel: Stacked 2D histogram of $\varepsilon$ vs. galactocentric radius, $r$, where the colour scale indicates the joint weighted density, $\mathrm{W}(\varepsilon,r)/\Sigma\mathrm{W}$, summed across all galaxies. The vertical subpanel to the right shows the normalised marginal (1D) distribution, $\mathrm{W}(\varepsilon)/\Sigma\mathrm{W}$. Right panel: Same as left panel, but for the $\delta$ parameter. Cyan dash-dotted lines in the main panel and black dash-dotted lines in the side panel denote the parameter range associated with $\leq 10$ per cent error in circular velocity inference based on the corresponding dynamical model. In the main subpanel, white solid and white dashed lines show the median and central $50$ per cent envelopes of the respective conditional distributions, $\mathrm{W}(\varepsilon|r)$ and $\mathrm{W}(\delta|r)$.
@@
Fig. 2 shows the radial distribution of both diagnostic parameters across the galaxy sample. The left panel corresponds to the mass-weighted density of $\varepsilon(r,\phi)$ – i.e. $\mathrm{W}(\varepsilon,r)/\Sigma\mathrm{W}$ – while the right panel shows the same for $\delta(r,\phi)$. In each case, the main subpanel presents a stacked 2D histogram of parameter values as a function of galactocentric radius, with colour indicating the joint distribution normalised by the total counts. Each galaxy here contributes 180 evenly-spaced azimuthal measurements per radial bin up to the radius marking the extent of its gaseous disc – $R_{\rm ext,H\textsc{i}}$, as defined in Sec. 2.2, enabling robust population-level statistics. The vertical subpanels display the normalised 1D distribution of the corresponding parameter. Cyan (main panels) and black (side panels) dash-dotted lines indicate the range of parameter values associated with $\leq 10$ per cent error in circular velocity inference under the relevant fluid model (i.e. $|\varepsilon|\lesssim 0.2$, $|\delta|\leq 0.1$). In the main subpanel, white solid and white dashed lines indicate the median and central $50^{\rm th}$ percentile of the respective conditional distributions, $\mathrm{W}(\varepsilon|r)$ and $\mathrm{W}(\delta|r)$. Several results emerge.
@@
As evident in the main panels, the vast majority of the samples contributing to these broad distributions originate from the inner disc regions ($r\lesssim 5$ kpc). This finding runs counter to expectations based on dynamical timescales, which predict more rapid relaxation in the central regions; instead, we find clear evidence for persistent perturbations. This is particularly relevant for core–cusp studies, since this radial range provides the strongest leverage for constraining dark matter density profiles. This is not the only hint that the centres of dwarf galaxies are continuously perturbed out of equilibrium. In fact, a recent investigation of radial trends in the radial acceleration scaling relation (RAR; McGaugh et al. 2016) using the SPARC sample (Lelli et al., 2016) shows that the inner regions – defined as $r<2R_{d}$, where $R_{d}$ is the stellar disc scale radius – are much less tight (in terms of intrinsic scatter) than the outer disc (see Fig. S1 in Ren et al. 2019). This trend can be interpreted as empirical evidence that inner regions of galaxies are indeed the most dynamically disturbed.
@@
Moreover, the $\delta$ parameter reveals a population-wide tendency to underestimate the true circular velocity in the inner disc ($\delta<0$), implying a systematic bias toward inferring smaller inner velocities, hence ‘cored’ profiles. As the white solid and dashed lines indicate, this bias arises preferentially within $r\lesssim 5$ kpc, where the median radial trend of $\delta$ departs from zero and the conditional distributions become increasingly skewed toward negative values. Although the median fractional offsets remain relatively modest – of order 10 per cent at $r\approx 2$ kpc – the growing asymmetry of the distributions implies that a substantial fraction of regions systematically underestimate the true circular velocity; crucially, this behaviour is mirrored, not mitigated, when adopting the equilibrium-based model. However, the $\delta$ distribution is broad in the inner kiloparsecs, with a minority of regions exhibiting the opposite trend ($\delta>0$). This bidirectional scatter suggests that some of the observed diversity in dwarf galaxy rotation curves could arise from disequilibrium physics rather than intrinsic variations in the underlying dark matter distributions. In other words, the rotational structure in the inner regions of gas discs is a biased and high-variance tracer of the underlying gravitational potential even before projection effects, and the limited ability of kinematic models to accurately recover azimuthal speeds, come into play.
@@ 4.1.1 Stellar feedback
Stellar feedback is a well-motivated source of ISM disequilibrium, particularly in shallow gravitational potentials where momentum and energy injection from massive stars can rival the disc binding energy (e.g. Dekel and Silk, 1986; Stinson et al., 2006; Hopkins et al., 2012; Kim and Ostriker, 2017). Energy deposition from SNe, stellar winds, and radiation pressure drives expanding shells, turbulence, and large-scale pressure gradients, each of which can displace gas from local force balance for timescales exceeding the local dynamical time (Martizzi et al., 2015; Orr et al., 2020). Observations support this picture: nearby dwarfs often exhibit asymmetric Hi profiles (e.g. Walter et al., 2008; Stilp et al., 2013) and vertically extended warm gas layers (Boomsma et al., 2008; Bacchini et al., 2020), with significant dispersive motions consistent with feedback-driven disturbances. A striking observational example of extreme disruption is the supergiant Hi shell, or ‘superbubble’, in IC2574 (Brinks et al., 2003), which also coincides with one of the most pronounced apparent dark matter cores known in a dwarf galaxy.
In our analysis, feedback-affected gas is defined as material lying within a projected radius of $1~\mathrm{kpc}$ in the warped disc midplane around either young stars (ages $\tau_{\star}\lesssim 100~\mathrm{Myr}$) or gas particles that have recently received kinetic injection from CCSNe ($t_{\rm lb}\leq 50~\mathrm{Myr}$). Particle positions are first assigned to their nearest point on the deprojected disc surface, and the selection is then made using a two-dimensional circular mask of radius $1~\mathrm{kpc}$ (see Sec. 2.5 for details). This choice reflects both the geometry of the discs and the physical dispersion of feedback-driven material: although CCSNe impart an isotropic target kick of $v_{\rm kick}=50~\mathrm{km~s^{-1}}$ in the rest frame of the star particle (Chaikin et al., 2023; Schaye et al., 2025), the residual velocity of affected gas in the rotating disc frame is typically smaller ($\sim 20~\mathrm{km~s^{-1}}$), implying characteristic displacements of $\sim 1~\mathrm{kpc}$ over $\sim 50~\mathrm{Myr}$ (ignoring other forces). Furthermore, Hi-dominated dwarf galaxy discs have expected vertical scale heights of $\sim 1~\mathrm{kpc}$ (Banerjee et al., 2011), making this spatial cut a physically motivated and conservative choice for identifying feedback-influenced regions in a population-wide analysis.
@@
[…]
@@ 4.1.2 Overdense and self-gravitating gas clumps
Local gravitational collapse within discs provides a natural route to hydrodynamical disequilibrium. Dense, self-gravitating clumps – often arising from gravitational instability in the cold ISM (e.g. Noguchi, 1999; Dekel et al., 2009; Genzel et al., 2011; Goldbaum et al., 2015) – develop internal dynamics that strongly deviate from the background shear flow.
@@
The middle row of Fig. 3 illustrates their spatial and dynamical imprint. In the left panel, the clump tracer field delineates spatially localised knots and filaments, frequently embedded within spiral arms or ring-like overdensities. The corresponding $\varepsilon$ map (middle) reveals that these same sites are often among the most extreme disequilibrium regions, with $\varepsilon$ commonly positive and exceeding unity. The right panel quantifies this: the clump-tagged distribution is markedly broader than the global disc-wide one, with a mild shift of the peak toward $\varepsilon\gtrsim 0$ and extended wings on both sides, reflecting a mix of accelerated collapse ($\varepsilon<0$) and migration ($\varepsilon>0$). This contrasts with the distribution characteristic of feedback zones, typically biased more strongly on the negative side (Sec. 4.1.1). In idealised scenarios, a clump drives a locally symmetric, converging flow in its own rest frame, such that $\partial_{t}v_{r}^{\prime}<0$, where $v_{r}^{\prime}$ is the radial component measured away from the clump (e.g., Penston, 1969). In the Eulerian galactocentric frame, because the ambient gas density decreases with radius, flux from the denser inward side exceed those from the outward side, biasing $\partial_{t}v_{r}$ towards the positive end. This naturally produces the mild positive shift in the clump-selected distribution. At the same time, many dense clumps either form within, or evolve into, regions where young stars subsequently inject feedback; sampling these mixed environments reintroduces the strongly negative tail characteristic of feedback-driven gas. The joint effect is a broad, wing-dominated distribution with a slight offset toward $\varepsilon>0$, as seen in Fig. 3.
While some clumps form or reside within feedback-active environments, a substantial fraction lie in quiescent or outer disc regions lacking obvious stellar-driven perturbations. This is evident in Fig. 3: because the same representative galaxy appears in the first two rows, the stellar-feedback (top-row) and clump-selected (middle-row) contours can be directly compared. Many overdense regions fall well outside the feedback-affected zones, yet still exhibit extreme $|\varepsilon|$, demonstrating that clump-driven dynamics generate strong disequilibria independently of stellar feedback. Observational analogues – such as the massive clumps in $z\sim 2$ discs (e.g. Elmegreen et al., 2007; Genzel et al., 2011), or in the nearby WLM dwarf galaxy (Albers et al., 2019; Mondal et al., 2021) – likewise show that such structures can persist without contemporaneous feedback. We therefore treat self-gravitating clumps as a separate, dynamically distinct driver of local disequilibrium.